Circular No. 7734

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ew IN ANONYMOUS GALAXY G. M. Hurst, Basingstoke, England, reports the discovery of an ap-

parent supernova (mag 17) by T. Boles, Coddenham, England, on a 30-s unfiltered CCD frame (limiting mag 18.5) taken with a 0.36-m telescope on Oct. 9.133 UT in the course of the U.K. Nova/Supernova Patrol. SN 2001ew is located at α = 3^{h}13^{m}06^{s}.98, δ = +42^{o}26′15′′.1 (equinox 2000.0), which is 5′′ east of the nucleus of its host galaxy. M. Armstrong, Rolvenden, England, reports the following position end figures measured from a confirming unfiltered CCD image taken with a 0.36-m Schmidt-Cassegrain telescope on Oct. 16.163: 07^{s}.46, 14′′.9. Another image obtained by Boles on Oct. 16.057 suggests that the object may have brightened slightly. A master image by Boles taken on Aug. 28 did not show the new object to limiting mag 18.5. An image taken by Armstrong on 2000 Oct. 6 also shows nothing at the position of SN 2001ew.

SUPERNOVA 2001eo IN UGC 3963 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2001eo (cf. IAUC 7726), obtained by P. Berlind on Oct. 13.49 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph) at Mt. Hopkins, shows it to be a type-Ia supernova after maximum light. The spectral-feature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 5 ± 2 days after maximum light.

SUPERNOVA 2001et IN MCG –03-51-9 Matheson et al. add that a spectrum of SN 2001et (cf. IAUC 7731),

taken by Berlind as above on Oct. 14.10 UT, shows it to be a type-II supernova.

COMET P/2001 Q6 (NEAT)

Visual m_{1} estimates: Sept. 22.00 UT, 14.0 (K. Sárneczky, Ágasvŕ, Hungary, 0.44-m reflector); Oct. 8.81, 11.8 (R. J. Bouma, Groningen, The Netherlands, 0.31-m reflector); 13.64, 11.9 (K. Yoshimoto, Yamaguchi, Japan, 0.25-m reflector); 14.74, 11.1 (Yoshimoto); 15.85, 11.4 (Bouma).

2001 October 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green